OpenStudy (anonymous):

Factor completely, then place the answer in the proper location on the grid. 14a^2 - 68a + 48

OpenStudy (anonymous):

a=4

OpenStudy (solomonzelman):

What a=4??!!! The common factor is 2, then factor it.

OpenStudy (anonymous):

74=5

OpenStudy (solomonzelman):

omg !!

OpenStudy (anonymous):

\[14 a^2-68a+48=2\left( 7 a^2- 34 a+24 \right)\] 7*24=168 34=28+6 28*6=168 write -34 a=-(28+6) a and make factors.